Glass Houses Pty has found that the cost of fitting new sliding doors is normally distributed, with an average cost per door of R12 750, and a standard deviation of R1 095. A sample of 105 new windows is randomly selected. Assuming normal distribution: Determine the standard error for this sample.


1
Expert's answer
2021-10-25T03:13:35-0400
"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{1095}{\sqrt{105}}\approx106.861058"